Solve the equation x(2x – 3) + 4 = 0
How do i solve this?
Thank you for yoour help.

Answer by ankor@dixie-net.com(22740)

Solve the equation
x(2x – 3) + 4 = 0
2x^2 - 3x + 4 = 0
Use the discriminant to see if this does have real solutions d = - b^2-4ac
a=2; b=-3; c=4
d =(-3)^2 + 4*2*4
d = 9 - 4*2*4
d = 9 - 32
d = -23, less than 0, no real roots